OpenStudy (ledah):

In a 30°- 60°- 90° right triangle, the longer leg is 6 times the square root of 3 cm. How long are the other two sides of the triangle?

ganeshie8 (ganeshie8):

hint : side opposite to large angle will be longer

OpenStudy (compassionate):

A + B + C = 180 \[B = 6\sqrt{3cm}\]

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

which side is the given \(6\sqrt{3} \ cm ?\)

OpenStudy (ledah):

guess that it would be the longer leg.

ganeshie8 (ganeshie8):

thats right

OpenStudy (compassionate):